OpenStudy (anonymous):

MEDAL AND FAN!!! Below are two different functions, f(x) and g(x). What can be determined about their y-intercepts

OpenStudy (anonymous):

f(X)

OpenStudy (anonymous):

g(x) 2 1 4 3 6 5

OpenStudy (thehumantorch10):

Both function are representing straight lines. Y intercept of f(x) is -1. Easy one as it is give in y = mx + c form. where 'm' is the slope of the line and 'c' is the y-intercept. And g(x) can be found as 'g(x)=x+2'. Again here the same y= mx+c form. So y-intercept is 2 So the Answer is B) The function g(x) has a higher y-intercept.

OpenStudy (thehumantorch10):

Do you understand?

OpenStudy (anonymous):

What is the equation of a vertical line passing through the point (3, −5)

OpenStudy (anonymous):

answer choices y=-5 x=-5 y=3 x=3

OpenStudy (thehumantorch10):

equation of horizontal line is y=-5 equation for vertical line is x=-3

OpenStudy (anonymous):

u mean x=3

OpenStudy (thehumantorch10):

oops yeah

OpenStudy (anonymous):

Which function below is the inverse of f(x) = x2 − 9

OpenStudy (thehumantorch10):

if x^2 = 9 then wat is x?

OpenStudy (anonymous):

plus and minus the square root of x+9 x^2 divide by 9 plus or minus 3 square root 3 1 divide by x^2 minus 9

OpenStudy (thehumantorch10):

jk

OpenStudy (thehumantorch10):

there you go bud

OpenStudy (anonymous):

it is d